# Analyzing Cognitive Differences Among Large Language Models through the Lens of Social Worldview

### v1

@@ F.1.4 Results
Appendix G Details for Paired T-test Results Across Different Prompt Settings (RQ2+RQ3)
Model t p Mean Diff. CI Lower CI Upper Significant Higher Setting 1 Setting 2
deepseek-chat 1.4187 0.1579 0.0250 -0.0098 0.0598 No Base Base Self-Awareness
1.9131 0.0575 0.0375 -0.0012 0.0762 No Self-Awareness Self-Awareness Feedback-None
0.8156 0.4159 0.0125 -0.0178 0.0428 No Self-Awareness Self-Awareness Feedback-Little
-1.4187 0.1579 -0.0125 -0.0299 0.0049 No Feedback-Most Self-Awareness Feedback-Most
gemini-2.0-flash -1.0429 0.2986 -0.0312 -0.0904 0.0279 No Self-Awareness Base Self-Awareness
8.7457 0.0000 0.4437 0.3435 0.5440 Yes Self-Awareness Self-Awareness Feedback-None
5.1478 0.0000 0.2500 0.1541 0.3459 Yes Self-Awareness Self-Awareness Feedback-Little
-1.7430 0.0833 -0.0187 -0.0400 0.0025 No Feedback-Most Self-Awareness Feedback-Most
gemma-3-27b-it 2.1339 0.0344 0.0437 0.0033 0.0842 Yes Base Base Self-Awareness
17.0341 0.0000 0.7625 0.6741 0.8509 Yes Self-Awareness Self-Awareness Feedback-None
11.1827 0.0000 0.4750 0.3911 0.5589 Yes Self-Awareness Self-Awareness Feedback-Little
-4.5149 $0$0.1250 -0.1797 -0.0703 Yes Feedback-Most Self-Awareness Feedback-Most
glm-4-9b-chat 1.9131 0.0575 0.0375 -0.0012 0.0762 No Base Base Self-Awareness
-8.9348 $0$0.4062 -0.4960 -0.3165 Yes Feedback-None Self-Awareness Feedback-None
-10.2261 $0$0.4313 -0.5145 -0.3480 Yes Feedback-Little Self-Awareness Feedback-Little
-11.0091 $0$0.4562 -0.5381 -0.3744 Yes Feedback-Most Self-Awareness Feedback-Most
gpt-4o 0.2418 0.8092 0.0063 -0.0448 0.0573 No Base Base Self-Awareness
4.5754 0.0000 0.1375 0.0781 0.1969 Yes Self-Awareness Self-Awareness Feedback-None
2.8928 0.0044 0.0500 0.0159 0.0841 Yes Self-Awareness Self-Awareness Feedback-Little
0.0000 1.0000 0.0000 0.0000 0.0000 No Self-Awareness Feedback-Most
internlm2_5-20b-chat -2.3415 0.0204 -$0$0.0922 -0.0078 Yes Self-Awareness Base Self-Awareness
5.4164 0.0000 0.2250 0.1430 0.3070 Yes Self-Awareness Self-Awareness Feedback-None
2.7059 0.0076 0.1125 0.0304 0.1946 Yes Self-Awareness Self-Awareness Feedback-Little
-5.0347 $0$0.1375 -0.1914 -0.0836 Yes Feedback-Most Self-Awareness Feedback-Most
llama-3.3-70b-instruct 3.3781 0.0009 0.0938 0.0389 0.1486 Yes Base Base Self-Awareness
1.1559 0.2495 0.0250 -0.0177 0.0677 No Self-Awareness Self-Awareness Feedback-None
-1.0000 0.3188 -0.0187 -0.0558 0.0183 No Feedback-Little Self-Awareness Feedback-Little
-5.0467 $0$0.1688 -0.2348 -0.1027 Yes Feedback-Most Self-Awareness Feedback-Most
phi-4 -2.9557 0.0036 -0.0625 -0.1043 -0.0207 Yes Self-Awareness Base Self-Awareness
-3.2558 0.0014 -0.0625 -0.1004 -0.0246 Yes Feedback-None Self-Awareness Feedback-None
-3.5905 0.0004 -0.0750 -0.1163 -0.0337 Yes Feedback-Little Self-Awareness Feedback-Little
-5.0347 $0$0.1375 -0.1914 -0.0836 Yes Feedback-Most Self-Awareness Feedback-Most
Qwen2.5-72B-Instruct -2.6971 0.0077 -0.0437 -0.0758 -0.0117 Yes Self-Awareness Base Self-Awareness
1.7430 0.0833 0.0187 -0.0025 0.0400 No Self-Awareness Self-Awareness Feedback-None
0.0000 1.0000 0.0000 0.0000 0.0000 No Self-Awareness Feedback-Little
0.0000 1.0000 0.0000 0.0000 0.0000 No Self-Awareness Feedback-Most
Table 11: Paired T-Test: Prompt vs. Prompt (same LLM) Comparisons for Egalitarianism — Combined Across LLMs. A dash in the “Higher” column means the two settings scored exactly the same.
Model t p Mean Diff. CI Lower CI Upper Significant Higher Setting 1 Setting 2
deepseek-chat 2.6017 0.0102 0.0875 0.0211 0.1539 Yes Base Base Self-Awareness
-2.6017 0.0102 -0.0875 -0.1539 -0.0211 Yes Feedback-None Self-Awareness Feedback-None
-4.2226 $0$0.1375 -0.2018 -0.0732 Yes Feedback-Little Self-Awareness Feedback-Little
-6.0720 $0$$0$0.2651 -0.1349 Yes Feedback-Most Self-Awareness Feedback-Most
gemini-2.0-flash 0.3769 0.7067 0.0125 -0.0530 0.0780 No Base Base Self-Awareness
-4.7946 $0$$0$0.4236 -0.1764 Yes Feedback-None Self-Awareness Feedback-None
-6.2043 $0$0.3438 -0.4532 -0.2343 Yes Feedback-Little Self-Awareness Feedback-Little
-4.3477 $0$0.1062 -0.1545 -0.0580 Yes Feedback-Most Self-Awareness Feedback-Most
gemma-3-27b-it 1.1482 0.2526 0.0312 -0.0225 0.0850 No Base Base Self-Awareness
1.5427 0.1249 0.1187 -0.0333 0.2708 No Self-Awareness Self-Awareness Feedback-None
-2.1562 0.0326 -0.1250 -0.2395 -0.0105 Yes Feedback-Little Self-Awareness Feedback-Little
-4.2517 $0$0.1313 -0.1922 -0.0703 Yes Feedback-Most Self-Awareness Feedback-Most
glm-4-9b-chat -4.0556 0.0001 -0.0938 -0.1394 -0.0481 Yes Self-Awareness Base Self-Awareness
-12.9698 $0$0.7125 -0.8210 -0.6040 Yes Feedback-None Self-Awareness Feedback-None
-15.9715 $0$0.7750 -0.8708 -0.6792 Yes Feedback-Little Self-Awareness Feedback-Little
-21.1806 $0$0.8375 -0.9156 -0.7594 Yes Feedback-Most Self-Awareness Feedback-Most
gpt-4o 1.8739 0.0628 0.0688 -0.0037 0.1412 No Base Base Self-Awareness
0.4461 0.6561 0.0187 -0.0643 0.1018 No Self-Awareness Self-Awareness Feedback-None
0.8156 0.4159 0.0125 -0.0178 0.0428 No Self-Awareness Self-Awareness Feedback-Little
-2.0191 0.0452 -0.0250 -0.0495 -0.0005 Yes Feedback-Most Self-Awareness Feedback-Most
internlm2_5-20b-chat -1.7430 0.0833 -0.0187 -0.0400 0.0025 No Self-Awareness Base Self-Awareness
-4.3448 $0$0.2562 -0.3727 -0.1398 Yes Feedback-None Self-Awareness Feedback-None
-8.4771 $0$0.4562 -0.5625 -0.3500 Yes Feedback-Little Self-Awareness Feedback-Little
-27.9861 $0$0.8313 -0.8899 -0.7726 Yes Feedback-Most Self-Awareness Feedback-Most
llama-3.3-70b-instruct 2.9019 0.0042 0.1625 0.0519 0.2731 Yes Base Base Self-Awareness
3.2498 0.0014 0.1062 0.0417 0.1708 Yes Self-Awareness Self-Awareness Feedback-None
2.1625 0.0321 0.0500 0.0043 0.0957 Yes Self-Awareness Self-Awareness Feedback-Little
-0.5988 0.5502 -0.0187 -0.0806 0.0431 No Feedback-Most Self-Awareness Feedback-Most
phi-4 0.9040 0.3674 0.0187 -0.0222 0.0597 No Base Base Self-Awareness
-5.4378 $0$0.2562 -0.3493 -0.1632 Yes Feedback-None Self-Awareness Feedback-None
-5.6414 $0$0.2437 -0.3291 -0.1584 Yes Feedback-Little Self-Awareness Feedback-Little
-4.4387 $0$0.1313 -0.1896 -0.0729 Yes Feedback-Most Self-Awareness Feedback-Most
Qwen2.5-72B-Instruct 2.3915 0.0179 0.0625 0.0109 0.1141 Yes Base Base Self-Awareness
-3.8682 0.0002 -0.1062 -0.1605 -0.0520 Yes Feedback-None Self-Awareness Feedback-None
-5.4440 $0$0.1688 -0.2300 -0.1075 Yes Feedback-Little Self-Awareness Feedback-Little
-3.5905 0.0004 -0.0750 -0.1163 -0.0337 Yes Feedback-Most Self-Awareness Feedback-Most
Table 12: Paired T-Test: Prompt vs. Prompt (same LLM) Comparisons for Fatalism — Combined Across LLMs. A dash in the “Higher” column means the two settings scored exactly the same.
Model t p Mean Diff. CI Lower CI Upper Significant Higher Setting 1 Setting 2
deepseek-chat 1.7079 0.0896 0.0437 -0.0068 0.0943 No Base Base Self-Awareness
1.3908 0.1662 0.0312 -0.0131 0.0756 No Self-Awareness Self-Awareness Feedback-None
1.0000 0.3188 0.0125 -0.0122 0.0372 No Self-Awareness Self-Awareness Feedback-Little
-3.2558 0.0014 -0.0625 -0.1004 -0.0246 Yes Feedback-Most Self-Awareness Feedback-Most
gemini-2.0-flash -0.5988 0.5502 -0.0187 -0.0806 0.0431 No Self-Awareness Base Self-Awareness
6.8360 0.0000 0.4313 0.3067 0.5558 Yes Self-Awareness Self-Awareness Feedback-None
3.9391 0.0001 0.2625 0.1309 0.3941 Yes Self-Awareness Self-Awareness Feedback-Little
-4.4894 $0$0.1125 -0.1620 -0.0630 Yes Feedback-Most Self-Awareness Feedback-Most
gemma-3-27b-it -0.8312 0.4071 -0.0187 -0.0633 0.0258 No Self-Awareness Base Self-Awareness
10.5372 0.0000 0.6250 0.5079 0.7421 Yes Self-Awareness Self-Awareness Feedback-None
8.9708 0.0000 0.4250 0.3314 0.5186 Yes Self-Awareness Self-Awareness Feedback-Little
-2.5383 0.0121 -0.0563 -$0$0.0125 Yes Feedback-Most Self-Awareness Feedback-Most
glm-4-9b-chat 0.0000 1.0000 0.0000 -0.0392 0.0392 No Base Self-Awareness
-5.4918 $0$0.3125 -0.4249 -0.2001 Yes Feedback-None Self-Awareness Feedback-None
-6.6071 $0$0.3688 -0.4790 -0.2585 Yes Feedback-Little Self-Awareness Feedback-Little
-12.3598 $0$0.5250 -0.6089 -0.4411 Yes Feedback-Most Self-Awareness Feedback-Most
gpt-4o 1.7159 0.0881 $0$0.0075 0.1075 No Base Base Self-Awareness
6.8252 0.0000 0.2687 0.1910 0.3465 Yes Self-Awareness Self-Awareness Feedback-None
3.9386 0.0001 0.1000 0.0499 0.1501 Yes Self-Awareness Self-Awareness Feedback-Little
0.0000 1.0000 0.0000 0.0000 0.0000 No Self-Awareness Feedback-Most
internlm2_5-20b-chat -2.1339 0.0344 -0.0437 -0.0842 -0.0033 Yes Self-Awareness Base Self-Awareness
6.5388 0.0000 0.3375 0.2356 0.4394 Yes Self-Awareness Self-Awareness Feedback-None
2.0336 0.0437 0.1187 0.0034 0.2341 Yes Self-Awareness Self-Awareness Feedback-Little
-10.4302 $0$0.4062 -0.4832 -0.3293 Yes Feedback-Most Self-Awareness Feedback-Most
llama-3.3-70b-instruct 2.6484 0.0089 0.0813 0.0207 0.1418 Yes Base Base Self-Awareness
4.8435 0.0000 0.1500 0.0888 0.2112 Yes Self-Awareness Self-Awareness Feedback-None
2.5739 0.0110 0.0500 0.0116 0.0884 Yes Self-Awareness Self-Awareness Feedback-Little
-0.9281 0.3548 -0.0312 -0.0978 0.0353 No Feedback-Most Self-Awareness Feedback-Most
phi-4 -2.5673 0.0112 -0.0688 -0.1216 -0.0159 Yes Self-Awareness Base Self-Awareness
-5.4378 $0$0.2562 -0.3493 -0.1632 Yes Feedback-None Self-Awareness Feedback-None
-5.6414 $0$0.2437 -0.3291 -0.1584 Yes Feedback-Little Self-Awareness Feedback-Little
-3.4261 0.0008 -0.0688 -0.1084 -0.0291 Yes Feedback-Most Self-Awareness Feedback-Most
Qwen2.5-72B-Instruct 0.6313 0.5288 0.0125 -0.0266 0.0516 No Base Base Self-Awareness
1.7430 0.0833 0.0187 -0.0025 0.0400 No Self-Awareness Self-Awareness Feedback-None
-0.8156 0.4159 -0.0125 -0.0428 0.0178 No Feedback-Little Self-Awareness Feedback-Little
NaN nan 0.0000 0.0000 0.0000 No Feedback-Most Self-Awareness Feedback-Most
Table 13: Paired T-Test: Prompt vs. Prompt (same LLM) Comparisons for Hierarchy — Combined Across LLMs. A dash in the “Higher” column means the two settings scored exactly the same.
Model t p Mean Diff. CI Lower CI Upper Significant Higher Setting1 Setting2
deepseek-chat 1.1559 0.2495 0.0250 -0.0177 0.0677 No Base Base Self-Awareness
4.4387 0.0000 0.1313 0.0729 0.1896 Yes Self-Awareness Self-Awareness Feedback-None
1.2937 0.1976 0.0312 -0.0165 0.0790 No Self-Awareness Self-Awareness Feedback-Little
-3.7498 0.0002 -0.0813 -0.1240 -0.0385 Yes Feedback-Most Self-Awareness Feedback-Most
gemini-2.0-flash -1.1482 0.2526 -0.0312 -0.0850 0.0225 No Self-Awareness Base Self-Awareness
11.7443 0.0000 0.6000 0.4991 0.7009 Yes Self-Awareness Self-Awareness Feedback-None
9.2333 0.0000 0.4125 0.3243 0.5007 Yes Self-Awareness Self-Awareness Feedback-Little
-1.7430 0.0833 -0.0187 -0.0400 0.0025 No Feedback-Most Self-Awareness Feedback-Most
gemma-3-27b-it -0.8939 0.3727 -0.0250 -0.0802 0.0302 No Self-Awareness Base Self-Awareness
24.8364 0.0000 1.0875 1.0010 1.1740 Yes Self-Awareness Self-Awareness Feedback-None
19.4056 0.0000 0.7500 0.6737 0.8263 Yes Self-Awareness Self-Awareness Feedback-Little
-4.5754 $0$0.1375 -0.1969 -0.0781 Yes Feedback-Most Self-Awareness Feedback-Most
glm-4-9b-chat 1.6761 0.0957 0.0312 -0.0056 0.0681 No Base Base Self-Awareness
-5.6116 $0$0.2875 -0.3887 -0.1863 Yes Feedback-None Self-Awareness Feedback-None
-7.8385 $0$0.3937 -0.4930 -0.2945 Yes Feedback-Little Self-Awareness Feedback-Little
-14.1254 $0$0.5687 -0.6483 -0.4892 Yes Feedback-Most Self-Awareness Feedback-Most
gpt-4o -0.9281 0.3548 -0.0312 -0.0978 0.0353 No Self-Awareness Base Self-Awareness
8.4125 0.0000 0.3312 0.2535 0.4090 Yes Self-Awareness Self-Awareness Feedback-None
4.2032 0.0000 0.1000 0.0530 0.1470 Yes Self-Awareness Self-Awareness Feedback-Little
0.0000 1.0000 0.0000 0.0000 0.0000 No Self-Awareness Feedback-Most
internlm2_5-20b-chat -1.6116 0.1090 -0.0375 -0.0835 0.0085 No Self-Awareness Base Self-Awareness
8.4359 0.0000 0.3625 0.2776 0.4474 Yes Self-Awareness Self-Awareness Feedback-None
5.3440 0.0000 0.2562 0.1615 0.3510 Yes Self-Awareness Self-Awareness Feedback-Little
-6.5502 $0$0.2125 -0.2766 -0.1484 Yes Feedback-Most Self-Awareness Feedback-Most
llama-3.3-70b-instruct 4.0189 0.0001 0.1812 0.0922 0.2703 Yes Base Base Self-Awareness
1.1349 0.2581 0.0187 -0.0139 0.0514 No Self-Awareness Self-Awareness Feedback-None
0.5761 0.5653 0.0063 -0.0152 0.0277 No Self-Awareness Self-Awareness Feedback-Little
-3.2455 0.0014 -0.0813 -0.1307 -0.0318 Yes Feedback-Most Self-Awareness Feedback-Most
phi-4 -2.0863 0.0386 -0.0563 -0.1095 -0.0030 Yes Self-Awareness Base Self-Awareness
-4.3749 $0$0.1187 -0.1724 -0.0651 Yes Feedback-None Self-Awareness Feedback-None
-4.3749 $0$0.1187 -0.1724 -0.0651 Yes Feedback-Little Self-Awareness Feedback-Little
-4.9012 $0$0.1313 -0.1841 -0.0784 Yes Feedback-Most Self-Awareness Feedback-Most
Qwen2.5-72B-Instruct -1.6761 0.0957 -0.0312 -0.0681 0.0056 No Self-Awareness Base Self-Awareness
1.4187 0.1579 0.0125 -0.0049 0.0299 No Self-Awareness Self-Awareness Feedback-None
0.0000 1.0000 0.0000 0.0000 0.0000 No Self-Awareness Feedback-Little
0.0000 1.0000 0.0000 0.0000 0.0000 No Self-Awareness Feedback-Most
Table 14: Paired T-Test: Prompt vs. Prompt (same LLM) Comparisons for Individualism — Combined Across LLMs. A dash in the “Higher” column means the two settings scored exactly the same.
Appendix H Details for Evaluating Social Feedback Effects Analysis (RQ3)
H.1 Statistical Methodology
To quantify the influence of social feedback on the cognitive attitudes of Large Language Models (LLMs), we developed a two-fold statistical framework comprising a repeated-measures analysis of variance (rm-ANOVA) with Bonferroni-corrected pairwise comparisons (RQ3a), and a polynomial trend analysis evaluating the dose-response relationship between varying feedback intensities and LLM cognitive outcomes (RQ3b).
Repeated-Measures ANOVA and Post-hoc Analysis (RQ3a).
We first employed a repeated-measures ANOVA to identify differences across experimental conditions, specifically comparing the Social Self-Awareness condition to various intensities of social feedback (Feedback-None, Feedback-Less, and Feedback-Most). Let $Y_{ijk}$ represent the observed cognitive score of the $i$-th LLM ($i=1,\dots,N$) measured along the $j$-th cognitive dimension (e.g., Egalitarianism, Individualism; $j=1,\dots,J$) under the $k$-th experimental condition, where

### v2

@@ D.1.4 Results
Appendix E Details for Paired T-test Results (Study 2)
Table 12: Paired T-Test: Self-Awareness vs. Base Comparisons for Egalitarianism — Combined Across LLMs. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
Model $t$ $p$ Mean Diff. CI Lower CI Upper Comparison
deepseek-chat -1.4187 0.1579 -0.0250 -0.0598 0.0098 Self-Awareness < Base
gemini-2.0-flash 1.0429 0.2986 0.0312 -0.0279 0.0904 Self-Awareness > Base
gemma-3-27b-it -2.1339 0.0344 -0.0437 -0.0842 -0.0033 Self-Awareness < Base *
glm-4-9b-chat -1.9131 0.0575 -0.0375 -0.0762 0.0012 Self-Awareness < Base
gpt-4o -0.2418 0.8092 -0.0063 -0.0573 0.0448 Self-Awareness < Base
internlm2_5-20b-chat 2.3415 0.0204 0.0500 0.0078 0.0922 Self-Awareness > Base *
llama-3.3-70b-instruct -3.3781 0.0009 -0.0938 -0.1486 -0.0389 Self-Awareness < Base ***
phi-4 2.9557 0.0036 0.0625 0.0207 0.1043 Self-Awareness > Base **
Qwen2.5-72B-Instruct 2.6971 0.0077 0.0437 0.0117 0.0758 Self-Awareness > Base **
Table 13: Paired T-Test: Self-Awareness vs. Base Comparisons for Fatalism — Combined Across LLMs. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
Model $t$ $p$ Mean Diff. CI Lower CI Upper Comparison
deepseek-chat -2.6017 0.0102 -0.0875 -0.1539 -0.0211 Self-Awareness < Base *
gemini-2.0-flash -0.3769 0.7067 -0.0125 -0.0780 0.0530 Self-Awareness < Base
gemma-3-27b-it -1.1482 0.2526 -0.0312 -0.0850 0.0225 Self-Awareness < Base
glm-4-9b-chat 4.0556 0.0001 0.0938 0.0481 0.1394 Self-Awareness > Base ***
gpt-4o -1.8739 0.0628 -0.0688 -0.1412 0.0037 Self-Awareness < Base
internlm2_5-20b-chat 1.7430 0.0833 0.0187 -0.0025 0.0400 Self-Awareness > Base
llama-3.3-70b-instruct -2.9019 0.0042 -0.1625 -0.2731 -0.0519 Self-Awareness < Base **
phi-4 -0.9040 0.3674 -0.0187 -0.0597 0.0222 Self-Awareness < Base
Qwen2.5-72B-Instruct -2.3915 0.0179 -0.0625 -0.1141 -0.0109 Self-Awareness < Base *
Table 14: Paired T-Test: Self-Awareness vs. Base Comparisons for Hierarchy — Combined Across LLMs. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
Model $t$ $p$ Mean Diff. CI Lower CI Upper Comparison
deepseek-chat -1.7079 0.0896 -0.0437 -0.0943 0.0068 Self-Awareness < Base
gemini-2.0-flash 0.5988 0.5502 0.0187 -0.0431 0.0806 Self-Awareness > Base
gemma-3-27b-it 0.8312 0.4071 0.0187 -0.0258 0.0633 Self-Awareness > Base
glm-4-9b-chat
gpt-4o -1.7159 0.0881 -0.0500 -0.1075 0.0075 Self-Awareness < Base
internlm2_5-20b-chat 2.1339 0.0344 0.0437 0.0033 0.0842 Self-Awareness > Base *
llama-3.3-70b-instruct -2.6484 0.0089 -0.0813 -0.1418 -0.0207 Self-Awareness < Base **
phi-4 2.5673 0.0112 0.0688 0.0159 0.1216 Self-Awareness > Base *
Qwen2.5-72B-Instruct -0.6313 0.5288 -0.0125 -0.0516 0.0266 Self-Awareness < Base
Table 15: Paired T-Test: Self-Awareness vs. Base Comparisons for Individualism — Combined Across LLMs. Significance levels: * $p<.05$, ** $p<.01$, *** $p<.001$.
Model $t$ $p$ Mean Diff. CI Lower CI Upper Comparison
deepseek-chat -1.1559 0.2495 -0.0250 -0.0677 0.0177 Self-Awareness < Base
gemini-2.0-flash 1.1482 0.2526 0.0312 -0.0225 0.0850 Self-Awareness > Base
gemma-3-27b-it 0.8939 0.3727 0.0250 -0.0302 0.0802 Self-Awareness > Base
glm-4-9b-chat -1.6761 0.0957 -0.0312 -0.0681 0.0056 Self-Awareness < Base
gpt-4o 0.9281 0.3548 0.0312 -0.0353 0.0978 Self-Awareness > Base
internlm2_5-20b-chat 1.6116 0.1090 0.0375 -0.0085 0.0835 Self-Awareness > Base
llama-3.3-70b-instruct -4.0189 0.0001 -0.1812 -0.2703 -0.0922 Self-Awareness < Base ***
phi-4 2.0863 0.0386 0.0563 0.0030 0.1095 Self-Awareness > Base *
Qwen2.5-72B-Instruct 1.6761 0.0957 0.0312 -0.0056 0.0681 Self-Awareness > Base
Appendix F Details for Evaluating Social Feedback Effects Analysis (RQ3)
To quantify the impact of social feedback on the cognitive attitudes of Large Language Models (LLMs), we employed a repeated-measures analysis of variance (rm-ANOVA), accompanied by Bonferroni-corrected pairwise comparisons.
Repeated-Measures ANOVA and Post-hoc Analysis
We utilized repeated-measures ANOVA to assess differences across experimental conditions, specifically comparing the Social Self-Awareness condition with three feedback intensities: Feedback-None, Feedback-Less, and Feedback-Most. Let $Y_{ijk}$ represent the cognitive score of the $i$-th LLM ($i=1,\dots,N$), measured along the $j$-th cognitive dimension (e.g., Egalitarianism, Individualism; $j=1,\dots,J$) under the $k$-th experimental condition, where:
$k$ $\in\{\text{Awareness},\text{Feedback-None},$
$\quad\;\text{Feedback-Less},\text{Feedback-Most}\}.$
The repeated-measures ANOVA model is formulated as:
where $\mu$ denotes the grand mean, $\alpha_{j}$ captures the main effect of cognitive dimension, $\beta_{k}$ represents the main effect of experimental condition, $(\alpha\beta)_{jk}$ denotes the dimension-condition interaction, $s_{i}\sim N(0,\sigma_{s}^{2})$ represents random effects capturing LLM-specific variability, and $\varepsilon_{ijk}\sim N(0,\sigma^{2})$ is the residual error.
The global null hypothesis tested was:
$\begin{split}H_{0}:\quad\beta_{\text{Awareness}}&=\beta_{\text{Feedback-None}}\\ &=\beta_{\text{Feedback-Less}}\\ &=\beta_{\text{Feedback-Most}}.\end{split}$
Upon rejecting this null hypothesis at significance level $p<0.05$, we conducted Bonferroni-corrected pairwise post-hoc comparisons between the Social Self-Awareness condition and each feedback intensity. The paired differences for each dimension $j$ and feedback condition were calculated as:
$D_{ij}^{(\text{Awareness–Feedback})}=Y_{ij,\text{Awareness}}-Y_{ij,\text{Feedback}},$
and evaluated using paired $t$-statistics:
$t_{j}$ $=\frac{\bar{D}_{j}}{s_{D_{j}}/\sqrt{N}},$
$\text{where}\quad\bar{D}_{j}$ $=\frac{1}{N}\sum_{i=1}^{N}D_{ij}^{(\text{Awareness–Feedback})},$
$s_{D_{j}}$ $=\sqrt{\frac{\sum_{i=1}^{N}(D_{ij}^{(\text{Awareness–Feedback})}-\bar{D}_{j})^{2}}{N-1}}.$
Practical significance was quantified using partial eta-squared ($\eta_{p}^{2}$), calculated from ANOVA results as:
and standardized mean difference (Hedges’ $g$) for each post-hoc test:
F.1 Results
Table 16: Significant Bonferroni‑adjusted post‑hoc comparisons ($p_{\text{corr}}\leq.05$). Positive $T$ means the first condition in the pair outscored the second.
Model Dim. Pair Compared Conditions $\mathbf{T}$ $\mathbf{df}$ $p_{\text{corr}}$
Qwen2.5‑72B Fatal. SA–FL SA >> FL 6.90 19 $8.4\times 10^{-6}$
SA–FN SA >> FN 5.10 19 $3.8\times 10^{-4}$
DeepSeek‑V3 Hier. FL–FM FL << FM 3.27-3.27 19 .024
Indiv. FN–SA FN << SA 4.70-4.70 19 .00093
Gemini‑2.0 Egal. FM–FN FM >> FN 9.30 19 $1.0\times 10^{-7}$
Indiv. FM–FN FM >> FN 11.46 19 $3.4\times 10^{-9}$
Gemma‑27B Indiv. FN–SA FN << SA 53.10-53.10 19 $2.4\times 10^{-21}$
GLM‑9B Hier. SA–FM SA << FM 13.42-13.42 19 $2.3\times 10^{-10}$
GPT‑4o Hier. FN–SA FN << SA 7.13-7.13 19 $4.4\times 10^{-6}$
InternLM‑20B Fatal. FM–SA FM >> SA 39.91 19 $5.2\times 10^{-19}$
Table 17: Repeated‑measures ANOVA (condition main effect) for every model–dimension pair. The FF‑ratio, uncorrected $p$ value, and partial $\eta^{2}$ are reported. Grey rows indicate clearly non‑significant results ($p>0.30$). A dagger ( $\dagger$ ) marks results that survive Bonferroni correction at $p<.05$.
Model Dimension $\mathbf{F}$ $\mathbf{p}$ $\boldsymbol{\eta^{2}_{p}}$
Qwen2.5‑72B Egalitarianism 3.35 .0250 .150 $\dagger$
Fatalism 23.02 $6.9\times 10^{-10}$ .548 $\dagger$